### v3

# Modelling the Spread of New Information on Social Networks

Abstract
There has been considerable interest in modelling the spread of information on social networks using machine learning models. Here, we consider the problem of predicting the spread of new information, i.e. when a user propagates information about a topic previously unseen by the user. In existing work, information and users are randomly assigned to a test or training set, ensuring that both sets are drawn from the same distribution. In the spread of new information, the problem becomes an out-of-distribution generalisation classification task.
Our experimental results reveal that while existing algorithms, which predominantly use features derived from the content of messages, perform well when the training and test distributions are the same, these algorithms perform much worse when the test set is out-of-distribution, i.e. when the topic (hashtag) of the testing data is absent from the training data. We then show that if the message features are supplemented or replaced with features derived from users’ profile and past behaviour, the out-of-distribution prediction is greatly improved, with the F1 score increasing from 0.117 to 0.705. Our experimental results suggest that a significant component of reposting behaviour for previously unseen topics can be predicted from users’ profile and past behaviour, and is largely content-agnostic.
Keywords Online Social Media $\cdot$ Social Content Sharing $\cdot$ Repost/Retweet $\cdot$ Out-of-Distribution Generalisation $\cdot$ Prediction $\cdot$ Machine Learning $\cdot$ X (Twitter)
1 Introduction
Understanding and predicting user reposting behaviour on social media platforms is an area of interest across disciplines such as computer science, social science, political science, and marketing. This interest stems from the influence that individual behaviours have on information diffusion [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12], including how information is spread during marketing and political campaigns [4, 5, 6, 7, 5, 10], and how misinformation and rumours spread [9, 8, 7, 10, 12].
The prediction of reposting has been formulated as a supervised classification task. The objective is to train a binary classifier to estimate the likelihood of a recipient reposting a message from a sender. In this setting, the dataset is typically divided into training and test sets through random assignment, ensuring that both sets are drawn from the same underlying data distribution for fair and stable evaluation.
On social media platforms, new topics appear regularly, e.g. those related to breaking news. Thus, in addition to the typical in-distribution classification, out-of-distribution prediction is also needed, where a model is trained on existing data related to previous topics (defined by hashtags) and is then used to predict a user’s reaction to a message related to a new, previously unseen, topic. Such a model can be useful for many applications, and it is more challenging, as it involves handling a distributional shift between user groups of known and unknown topics.
In this paper, we first review previous work on predicting repostings on X (formerly Twitter) [13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 3, 37, 38, 39, 2, 40, 41, 42, 1, 43, 44, 45, 46, 47, 48]. Existing algorithms predominantly focused on the textual data, including content of reposted messages and user historical posts. We argue that features of the sender and the recipient are also relevant to the reposting of a message. We study two types of data: (i) the message data, which contain the textual content of messages under consideration for reposting, (ii) the user data, which contain the sender and recipient’s profiles, their following lists, and their historical posts encoding their past activities, interests, and interactions with other users. We identified a total of 303 features from these data, including 78 message-related features and 225 user-related features. We consider three classification models: a decision tree model, a neural network model, and the BERT model [49], which is a widely studied Large Language Model (LLM). Using data collected on X, we conduct experiments for both in-distribution and out-of-distribution predictions.
Our results show that while the prediction models, using only message-related features, perform well for in-distribution prediction, they perform much worse for out-of-distribution prediction. When the message-related features are supplemented or replaced with user-related features, the out-of-distribution performance greatly improves with the F1 score increasing from 0.117 to 0.705.
Our work highlights the importance of user-related features for reposting prediction—features that are currently underutilized. Interestingly, our experimental results suggest that a significant component of reposting behaviour for previously unseen topics is largely content-agnostic. This work provides novel insights into the mechanisms underlying information spreading on social media.
2 Background
2.1 Reposting on Social Media Platforms
X (formerly Twitter) is one of the most popular social media platforms. It allows users to ‘post’ short messages (also called tweets), follow other users to curate their content feed, and interact with other users through actions, such as ‘like’ and ‘repost’. When a user posts a message, the message is shown to the user’s followers as well as other users who might be interested as determined by X’s recommendation algorithm.
A user can repostrepost (or ‘retweet’) a message. The act of reposting is often considered as an endorsement of a message and/or its creator. Reposting plays an important role in the diffusion of information on social media. On X, a user can also ‘quote’ or ‘reply’ to a message. Similarly to a repost, these actions will add the message to the timelines of the user’s followers and, therefore, will have an equivalent function as the repost in terms of information diffusion. In our X data, reposts account for 80% of the messages, and quotes and replies account for 2%. For simplicity, in this study we merge quotes and replies into reposts. A user can also ‘like’ a message to signal their endorsement, but this action does not add the message to the timelines of the user’s followers and has limited impact on information diffusion. Thus, the action of ‘like’ is not considered in this study.
Facebook and Sina Weibo also have similar repost functions (referred to as ‘share’ on Facebook). Despite the availability of datasets from these platforms, most research in literature has primarily focused on X due to its widespread use and greater accessibility until recently. Although Instagram has a ‘share’ function and TikTok offers a ‘repost’ feature, these differ from the ‘repost’ action discussed in this study, as they typically involve private forwarding of messages or posts to specific users or group chats.
2.2 Reposting Prediction
In this study, reposting prediction is defined as to predict whether a recipient will repost a message received from a sender. A reposting action involves three elements: (1) a message (M); (2) a sender ($U_{S}$) of the message who creates the message; and (3) a recipient ($U_{R}$) of the message via following or recommendation.
Specifically, we want to learn a binary classification model $f$, to predict the label $y$,
$f:\{M,U_{S},U_{R}\}\rightarrow y\in\{0,1\}\,$, (1)
where $y=1$ means the recipient will repost the message received from the sender, and $y=0$ otherwise.
Table 1: Message and User Data and Features.
Data Feature Type Feature Sub-type Example feature Note
(# of features) (# of features)
Message Data Message-Related Features
Data(M) [0.1cm] Text content [0.1cm] of message M (78) Topic (39) M_TopicM5 Likelihood of a message related to topic ‘family’.
Language (10) M_Grammar1 Grammatical correctness of a message.
Readability (11) M_Readability1 A measure of difficulty in understanding a message.
Sentiment (5) M_Sentiment1 Negative sentiment score of a message.
Emotion (8) M_Emotion1 Probability of a message expressing emotion ‘anger’.
Hate-Speech (4) M_Hate1 A measure of aggressiveness of a message.
Hashtag (1) M_Hashtag The presence of a specific hashtag in a message.
User Data User-Related Features for the Sender $U_{S}$ and the Recipient $U_{R}$
Data(U-P) [0.1cm] Profile and [0.1cm] friend list of [0.1cm] $U_{S}$ and $U_{R}$ U-P (30) Profile of $U_{S}$ (12) U-P_S_FollowerNum The number of followers of the Sender $U_{S}$.
Profile of $U_{R}$ (12) U-P_R_FollowerNum The number of followers of the Recipient $U_{R}$.
Network of $U_{S}$ (3) U-P_S_LeaderRank A measure of the Sender $U_{S}$’s influence.
Network of $U_{R}$ (3) U-P_R_LeaderRank A measure of the Recipient $U_{R}$’s influence.
Data(U-HA) [0.1cm] Public metrics [0.1cm] of historical [0.08cm] messages of [0.05cm] $U_{S}$ and $U_{R}$ U-HA (38) Popularity of $U_{S}$ (4) U-HA_S_LikedRate Avg. # of ‘Likes’ received by latest 50 messages of $U_{S}$.
Popularity of $U_{R}$(4) U-HA_R_LikedRate Avg. # of ‘Likes’ received by latest 50 messages of $U_{R}$.
Activity of $U_{S}$ (7) U-HA_S_InteractivePer Percentage of reposts in $U_{S}$’s latest 50 messages.
Activity of $U_{R}$(7) U-HA_R_InteractivePer Percentage of reposts in $U_{R}$’s latest 50 messages.
Interaction (16) U-HA_RS_Mention Number of $U_{R}$’s latest 50 messages that mentioned $U_{S}$.
Data(U-HM) U-HM (157) M features of $U_{S}$ (78) U-HM_S_Grammar1 (Each M feature is averaged over $U_{S}$’s latest 50 messages.)
Text content M features of $U_{R}$ (78) U-HM_R_Grammar1 (Each M feature is averaged over $U_{R}$’s latest 50 messages.)
of historical Topic Similarity (1) U-HM_SR_TopicSim Similarity between historical topics of 50 latest messages
messages of of $U_{S}$ and $U_{R}$.
$U_{S}$ and $U_{R}$
2.3 Existing Works
2.3.1 Related Works
Many works on the broader topic of information diffusion are related to the reposting action on social media. Some of them do not consider all the three elements defined in Equation (1), namely a message, a sender and a recipient. Some study messages only, some focus on messages and their senders only, and some consider messages and their recipients (i.e. reposters) only. For example, [26, 27] predicted the popularity of a message; [28, 29] predicted the incremental size of spreading tree for a given message over time; [30] predicted the time gap between reposts of a given message; [31] predicted the probability of a given message being reposted (by any user), and [32, 33, 34, 35, 36] predicted the next user who is going to repost a given message based on the sequence of previous reposters of the message and the social relations among these reposters.
2.3.2 Reposting Prediction
In this study, we focus on existing works on the reposting prediction task as defined in Equation (1) involving all the three elements.
These include a number of machine learning (ML) models, such as Support Vector Machines (SVM) [15, 16], Decision Trees  [15, 17, 13, 18], Logistic Regression  [19, 20], and factorisation models  [21, 22, 23, 24]. These models are based on features extracted from raw data for prediction. It is suggested in [15, 13] that the Decision Tree models often outperform other ML models.
Neural network (NN) models are also used for reposting prediction, including the Attention-based Convolutional Neural Network (SUA-ACNN) model [14] and the Bidirectional Long Short-Term Memory (Bi-LSTM) model [25]. NN models can learn directly from raw data and bypass the need for feature extraction.
2.3.3 Limitations of Previous Reposting Prediction
Reliance on Message-Related Data
Previous models primarily relied on data of messages, including messages being reposted and the historial posts of users. For example, the ML models typically used (a relatively small set of) features mostly extracted from the textual content of message data. The NN models used the message data exclusively. The Bi-LSTM model even ignored the User ID associated with a message. It is known, however, that user-related data, including users profile, their following list, and their past behaviour, interest and interactions, can have a significant influence on a user’s status, actions and decisions [50, 51, 16, 15]. For example, [16] suggested that a recipient is more likely to repost a message from a sender if the recipient follows the sender, and the two users have shared interests and prior interactions.
Lack of Out-of-Distribution Evaluation
Existing works typically evaluated model performance by splitting the data into training and test sets using random sampling [15, 16, 17, 13, 18, 19, 21, 22, 23, 24, 14, 25], where both sets contain data from the same set of topics. This setting, often referred to as in-distribution evaluation [52, 53], assumes that the training and test data are drawn from the same underlying distribution. However, different topics may associate with user groups with distinct feature distributions. None of prior works explicitly evaluated generalisation across different topics in the out-of-distribution setting, that is to train a model using data of some topics and then test the model using data of other topics, which can evaluate a model’s capacity in terms of predicting reposts associated with new, previously unseen information.
3 Method
3.1 X Data
We collected data from X using the former Twitter API with the Academic Access. We selected the 14 most popular hashtags in the trending list between July 27th and August 14th in 2022 and retrieved all available messages containing these hashtags during the same period.
We obtained 111,401 messages, each containing one of the 14 trending hashtags. From these messages, we identified 44,014 reposts (positive instances) that satisfied the following conditions: (1) the original message being reposted was also in our data and (2) the recipient created the repost within 24 hours since the original message was created by the sender. According to the study of half-life of X messages [54], users are unlikely to see a message more than 24 hours after its creation. Any reposting after 24 hours are mostly influenced by external factors, such as TV or other social media platforms.
These messages were posted by 79,707 unique users around the world. For each user, we collected their profile, following list, and up to 50 latest historical messages. A total of more than 3.8 million historical messages were collected. All users are anonymised in our data.
3.2 Message Data and Features
A message can have text and multimedia content, such as pictures, audio and video. Same as in prior works, in this study, we only consider the text content of a message.
In this study, the message data, Data(M), contain the text content of the 111,401 messages in our X data. From the message data, we extract 78 message-related features, M, which are used in prior works for reposting prediction.
These features characterise a message from different perspectives. As shown in Table 1, we divide them into sub-types, including topic  [43, 41, 40, 13, 17, 24], language style [44, 21, 42, 2], readability [42], sentiment [45, 3, 38, 17], emotion [1, 3, 37, 17], hate speech  [46], and hashtag. These features are calculated using publicly available language models and tools (see Appendix). In addition, the hashtag feature is recorded as a categorical label [21, 17].
3.3 User Data and Features
In this study, the user data, Data(U), contains the following three types of data related to a user.
  • Data(U-P): the profile of a user and the list of ‘friends’ that the user follows.
  • Data(U-HA): the metadata (or public metrics) of the latest 50 historical messages of a user, including (1) the type of a message, i.e. post/repost/quote/reply, (2) the creation time of a message, (3) the number of reposts/quotes/replies/likes a message has received, and (4) any usernames mentioned in a message. These metadata encode useful information of a user’s previous behaviours, activity patterns, and its interactions with other users.
  • Data(U-HM): the textual content of the latest 50 historical messages of a user, which characterise a user’s historical interests and communication style.
As shown in Table 1, we can extract three types of user-related features from the above three types of user data, respectively. More details of the features are provided in the Appendix.
3.3.1 User Profile Features (U-P)
These features include the number of followers/followees of a user [1, 40, 47, 2], the number of lists that a user is included in [16, 15], the number of messages a user has posted since its registration [16, 48, 40], whether a user is verified [21, 15], whether a URL is provided in a user’s profile [39], a user’s indegree [13] and LeaderRank [13], and whether a user is following another user [15].
3.3.2 User Historical Action Features (U-HA)
These features include, for example, the percentage of a user’s historical messages that are reposts [21, 13], the average time interval between a user’s posts [21], and the ratio of a sender’s historical messages that have mentioned a recipient [16, 15].
3.3.3 User Historical Message Features (U-HM)
These features include the aggregated message features (M) of a sender and those of a recipient, as well as an additional feature measuring topic similarity between the two users [13, 20].
3.4 Positive and Negative Instances
In this study, we define a positive instance as a reposting event $\{M(h),U_{S},U_{R},t\}$, where a recipient $U_{R}$, at time $t$, reposted a message $M(h)$ with a hashtag $h$, and the message was created by a sender $U_{S}$ within the 24 hours before $t$.
Correspondingly, a negative instance is $\{M^{*}(h),U^{*}_{S},U_{R},t\}$, where the same recipient $U_{R}$ did not repost another message $M^{*}(h)$ containing the same hashtag $h$, and the message $M^{*}(h)$ was created by sender $U^{*}_{S}$ (which can be the same as $U_{S}$) within the 24 hours before the same $t$. The messages in the positive and negative instances contained the same hashtag and were created in the same 24 hours, so that it is likely that $M^{*}(h)$ was also exposed to the recipient and therefore make the evaluation more relevant.
From the above X data, we created three datasets for reposting prediction with different ratios of positive to negative instances.
  • 1:1 dataset: For each positive instance, one negative instance is selected from our X data so that the message and the sender in the negative instance are most similar to those in the positive instance, i.e. the cosine distance between the feature vectors (containing all features of the message and the sender) of the positive instance and the chosen negative instance is the smallest among all negative instances.
  • 1:5 dataset: For each positive instance, the five most similar negative instances are selected, as defined above.
  • 1:10 dataset: For each positive instance, in addition to the 5 most similar negative instances, we add 5 negative instances randomly chosen from our X data.
Note that in all experiments in this study, if data is randomly split into training (and validation) and testing datasets, we remove any instance from the training dataset if the instance’s sender-recipient pair appears in the testing dataset. Such that, instances in the training and those in testing datasets always have different sender-recipient pairs. This is to prevent information leakage and ensure a fair and rigorous evaluation.
3.5 Prediction Models
3.5.1 Decision Tree (DT) Models
The DT models remain a competitive choice for reposting prediction [13, 15]. In this study, we will examine how the performance of DT models are affected when different types of features are used. We use the XGBoost decision tree model. We searched for the optimal hyper-parameters for each data split of each dataset. The most frequently used hyper-parameters include the maximum depth of 8, the learning rate of 0.3, the estimators of 100, the minimum child weight of 1, the subsample ratio of 1.0, and the positive class weight of 1.0. For more details, please refer to the Appendix.
3.5.2 Neural Network (NN) Models
Figure 1: Neural network (NN) models based on different types of data.
The NN models have become a popular choice for reposting prediction [14, 25]. NN models can learn directly from raw data and bypass the need for feature extraction.
As shown in Figure 1, the NN models used in this study are extended from the SUA-ACNN model [14]  [^1] The SUA-ACNN model only used textual data, but it used textual data from not only the message being considered for reposting, Data(M), but also historical messages of the sender and the recipient of the reposting, Data(U-HM). In our study, we separate message data from user-related data. by adding a component of Multi-Layer Perceptron (MLP)  [^2] The MLP component has three fully connected layers with 128, 128 and 64 units, respectively. The activation functions are set to ReLU. to incorporate additional user-related data, i.e. Data(U-P) and Data(U-HA). This allowed us to obtain a series of models using different types of data, including the NN-U model using the three user-related data, the NN-M model using the message data only, and the NN-ALL model using all types of data. NN models are trained with the Adam optimizer, the categorical cross-entropy loss, the learning rate of $0.001$, the batch size of $40$, and early stopping on validation loss.
3.5.3 BERT
We also study the BERT model [49], which is a widely studied Large Language Model (LLM). The model takes the message data as input and generates embeddings that represent the semantic meanings of the messages. The model then produced the prediction based on the embeddings. We configure BERT-base with the AdamW optimizer, the learning rate of 2e52e-5, the batch size of $32$, and the dropout rate of $0.1$, with early stopping on validation loss. Cross-entropy loss was used as the objective function.
3.6 Experiments
In this study, we conducted three experiments for reposting prediction using the above models. We examined whether the performance of reposting prediction can be improved by supplementing or replacing message-related data with user-related data. In addition, we evaluated the models for both in-distribution and out-of-distribution predictions.
3.6.1 F1 Score
The formula of F1 Score is shown below, where $TP$, FPFP, and $FN$ represent true positives, false positives, and false negatives, respectively.
$F1=\frac{TP}{TP+\frac{1}{2}(FP+FN)}$ (2)
3.6.2 Mean and Standard Deviation
For each prediction setting, we obtain ten F1 Scores. We calculate the mean $\boldsymbol{\mu}$ and standard deviation $\boldsymbol{\sigma}$ as defined below, where $j$ indexes the ten F1 Scores.
$\mu=\frac{1}{10}\sum_{j=1}^{10}F1_{j}$ (3)
$\sigma=\sqrt{\frac{1}{10}\sum_{j=1}^{10}(F1_{j}-\mu)^{2}}$ (4)
3.6.3 Overall Mean and Overall Standard Deviation
When predicting for an individual hashtag (in Experiments II and III), we obtain 10 F1 Scores and calculate the mean and the standard deviation as above, and then we further compute the overall mean $\boldsymbol{\bar{\mu}}$ and overall standard deviation $\boldsymbol{\bar{\sigma}}$ by treating results from the 14 hashtags as equally weighted components of a mixture distribution, defined as followings where $i$ indexes the 14 hashtags.
$\bar{\mu}=\frac{1}{14}\sum_{i=1}^{14}\mu_{i}$ (5)
$\bar{\sigma}=\sqrt{\frac{1}{14}\sum_{i=1}^{14}\sigma_{i}^{2}+\frac{1}{14}\sum_{i=1}^{14}(\mu_{i}-\bar{\mu})^{2}}$ (6)
4 Experiment I: In-Distribution Prediction For Mixed Hashtags
In this experiment, we mix data of all hashtags, and then performed Monte Carlo cross-validation 10 times, each randomly splitting the data 63/7/30 to the training, validation and test sets.
Table 2: In-Distribution Prediction for Mixed Hashtags
In-Distribution Training on 70% randomly selected data, and testing on the remaining 30% data.
Decision Tree (DT) model Neural Network (NN) model LLM model
Model DT-ALL DT-U DT-M NN-ALL NN-U NN-M BERT
Feature(ALL2) Feature(U3) Feature(M) Data(All) Data(U) Data(M) Data(M)
Dataset 1 F1 Score ($\mu$ $\pm$ $\sigma$4
[…]
1 Ratio of positive and negative instances. 2 ALL include both user-related features U and message-related features M. 3 U includes U-P, U-HA and U-HM. 4 The mean $\mu$ and standard deviation $\sigma$ are computed as in Eq. (3)-(4).
4.1 Results
As shown in Table 2, the DT-ALL model using all types of features (303 in total) consistently shows significantly improved performance than other models using fewer features across datasets with different ratios of positive and negative instances. Among models with reduced feature sets, the DT-U model (utilizing user-related features) outperforms the DT-M model (utilizing message-related features).
The NN-ALL model utilizing all types of data produces the strongest performance among NN models. The NN-M model and the BERT model both utilize only message-related data and they achieve similar performance. Notably, these two models are outperformed by the NN-U model which used only user-related data.
These results consistently indicate that the inclusion of user-related data is valuable for reposting prediction.
As expected, the performance of all models degrades as the proportion of negative instances increases in the dataset, which aligns with established understanding of classification challenges in imbalanced datasets. In the following of this paper, all results are based on the 1:5 dataset.
4.2 Comparison with Prior Works
In addition to the above results, we also conducted Experiment I using the decision tree model, TORS, in [13], which uses 37 features primarily derived from textual data of messages, and the neural network model, SUA-ACNN, in [14], which relies solely on textual data. On the 1:5 dataset, they achieved F1 scores of 0.737 and 0.739, respectively, which as expected are close to but lower than the DT-M model and the NN-M model.
Table 3: Top 20 Most Important Features based on Decision Tree Prediction.
Feature name Rank Feature Feature Import.
type sub-type value
U-P_R_FollowS 1 U-P network 0.229
U-HA_S_RetweetedRate 2 U-HA popularity 0.047
U-HA_S_LikedRate 3 U-HA popularity 0.045
U-HA_S_RepliedRate 4 U-HA popularity 0.042
U-HA_RS_MentionPer 5 U-HA interaction 0.015
M_TopicLDA7 1 6 M topic 0.013
U-HA_S_InteractivePer 7 U-HA activity 0.013
U-HA_S_TweetNum 8 U-HA activity 0.011
U-HA_S_RetweetPercent 9 U-HA activity 0.009
U-P_S_FollowR 10 U-P network 0.008
U-HA_RS_RepostLatency 11 U-HA interaction 0.008
U-HM_S_Readability7 12 U-HM readability 0.008
M_TopicLDA4 13 M topic 0.008
M_TopicLDA3 14 M topic 0.008
U-HM_S_TopicM7 15 U-HM topic 0.008
U-HM_S_TopicM11 16 U-HM topic 0.007
U-HM_S_TopicM19 17 U-HM topic 0.007
U-HA_RS_Mention 18 U-HA interaction 0.007
U-HA_S_TweetPercent 19 U-HA activity 0.007
U-P_S_Indegree 20 U-P network 0.007
Number of features in each type: U-HA: 10; U-HM: 4; U-P: 3; M: 3.
1 M_TopicLDA7 represents topics related to energy crisis, Brexit, and health issues. See full details of features in the Appendix.
4.3 Analysis of Feature Importance
To further analyse the relative importance of the 303 features, we obtained their importance values as a by-product of XGBoost decision tree prediction. The feature importance value represents the relative contribution of a feature to the decision tree model’s predictive performance. Specifically, it is computed as the reduction in the loss function contributed by each feature when splitting nodes, with larger reductions indicating higher importance. The feature importance values are normalised such that the sum for all features equals to one.
We obtained the feature importance values for all 303 features by training the DT-ALL model using all (instead of 70%) data in the 1:5 dataset. Table 3 shows the top 20 most important features. It is notable that the majority of top 20 features are user-related features, including 10 user historical action (U-HA) features, 4 user historical message (U-HM) features, and 3 user profile (U-P) features. The most important message (M) feature is merely ranked at the 6th position. This highlights the importance of user-related features for reposting predictions.
5 Experiment II: In-Distribution Prediction For Individual Hashtags
In this experiment, we conduct the in-distribution predictions for each of the 14 individual hashtags. We performed the same Monte Carlo cross-validation but on data of a single hashtag.
Table 5 presents the in-distribution prediction results for each individual hashtag. The results are broadly consistent with those observed from the mixed-hashtag setting, with minor variations across individual hashtags. Specifically, the overall ranking of models remains the same: for decision trees, DT-ALL outperforms DT-U, which in turn outperforms DT-M; for neural network models, NN-ALL achieves the highest performance, followed by NN-U, with NN-M and BERT producing similar lowest performance. These findings indicate that the relative effectiveness of user-related and message-related data is stable across different in-distribution conditions.
The splitting schemes can also impact results. For the two splitting schemes, the mixed hashtag scope, which leverages data across all 14 hashtags, generally produces better performance than the individual hashtag scope. This suggests benefits from increased diversity and volume of the training data.
Note that, for rigorous evaluation, we have also conducted additional temporal experiments where data of each individual hashtag are divided into time windows according to the timestamp of repost events (see additional temporal experiments in Appendix). The results and conclusions remain the same: models using user-related data consistently outperform those using message-related data.
Table 4: In-Distribution Prediction for Individual Hashtags on the 1:5 Dataset
In-Distribution For a given hashtag, training on 70% randomly selected data, test on remaining 30% data.
Decision Tree (DT) model Neural Network (NN) model LLM model
Model DT-ALL DT-U DT-M NN-ALL NN-U NN-M BERT
Feature(ALL) Feature(U) Feature(M) Data(All) Data(U) Data(M) Data(M)
Hashtag F1 Score ($\mu$ $\pm$ $\sigma$)
[…]
The mean, $\mu$, and standard deviation, $\sigma$, of F1 score for each hashtag are computed using Eq. (3)-(4). The overall mean, $\bar{\mu}$, and overall standard deviation, $\bar{\sigma}$, across 14 hashtags are computed using Eq. (5)-(6). Same in Table 5.
Table 5: Out-of-Distribution Prediction on the 1:5 Dataset
Out-of-distribution For a given hashtag, training on data of other hashtags, test on data of given hashtag.
Model DT-ALL DT-U DT-M NN-ALL NN-U NN-M BERT
Hashtag F1 Score ($\mu$ $\pm$ $\sigma$)
[…]
The expected F1 score for random guessing is 0.167, where positive and negative labels are randomly assigned to data at a ratio of 1:5.
6 Experiment III: Out-of-Distribution Prediction
We conducted out-of-distribution prediction for each of the 14 hashtags in our X data. For each given hashtag, we trained a model using data of the other 13 hashtags, and then tested the model with data of the given hashtag (the new, unseen topic). This split ensures an out-of-distribution condition, where the test messages come from a hashtag (topic) unseen during training  [^3] In our X data, we removed the small amount (less than 0.1%) of sender–recipient pairs that are overlapping across different hashtags. This ensures that sender–recipient pairs in the testing data are also unseen during training..
To obtain a stable and reliable estimate of model performance, for each given hashtag, we randomly divided the training data of the other 13 hashtags into 10 subsets, left one subset out for validation and used the other 9 subsets to train the model, and obtain prediction results. Then we repeated the process 10 times by leaving a different subset out each time. We evaluated the 10 trained models against the same testing data of the given hashtag under study, and then obtained the mean and standard deviation from ten F1 scores.
6.1 Results
Table 5 shows the out-of-distribution prediction by DT, NN and BERT models for each of the 14 hashtags on the 1:5 dataset. When based only on message-related data, all models show notably poor performance. Their overall F1 scores, $\bar{\mu}$, across 14 hashtags are as low as 0.117, 0.108 and 0.091, respectively, which are close to the expected F1 score of random guessing (0.167, see notes in Table 5). By comparison, the DT-ALL model and the NN-ALL model, using all types of data, perform significantly better.
Remarkably, the DT-U model and NN-U models, using only the user-related data, produce performance as good as, or even better than the DT-ALL and NN-ALL using all data. This suggests that user-related data are crucial for out-of-distribution prediction, where adding message-related data would have minimum benefit, or even negative impact, on performance.
The low performance of the NN-M model and the BERT model are unlikely due to any limitation in their architecture. Their underperformance is a result of exclusive reliance on message data. These two models learn from textual content of messages and therefore are bounded by the vocabularies present in the training data. This constrains their ability to generalise to messages of a new topic that are absent in the training data.
6.2 Discussion
6.2.1 In-Distribution vs. Out-of-Distribution
Figure 2 shows that the decision tree models’ performance for out-of-distribution prediction is always lower than that for in-distribution prediction. This is expected, as out-of-distribution prediction evaluates the model on a topic that is absent from the training set, whereas in-distribution prediction uses data of the same topics for both training and testing.
The gap between the two predictions is particularly pronounced for the model using message features M only, whose performance collapses for out-of-distribution prediction. This not only reveals the limited generalisation capability of models relying on message data, but also highlights the importance of conducting both in-distribution and out-of-distribution evaluations.
The model using U-HM features, extracted from the text content of users’ historical messages, performs similarly to message-based model (M). Its out-of-distribution performance declines noticeably, though not as severely as M. This suggests that, beyond M, U-HM also contributes little to out-of-distribution prediction compared with features of user profiles (U-P) and user historical actions (U-HA).
In contrast, three user-based models (U, U-P and U-HA) demonstrate significantly stronger predictive power, producing a performance as good as the model using ALL features for both in-distribution and out-of-distribution predictions. This is because these user-related features encode a user’s status and activity patterns that are independent from the latest trends on X and therefore can generalise across different topics.
6.2.2 The Decision Tree
Performance of the DT-ALL model, using all types of features, is higher that that of the NN-ALL model, using all types of data, in both in-distribution experiments (for mixed and individual hashtags) and out-of-distribution experiments.
This can perhaps be explained by the ability of decision trees to disregard uninformative features through their feature-splitting mechanism: at each node, the tree chooses the feature that best separates the classes, which not only reveals useful signals hidden in the data, but also effectively ignores weak or irrelevant features.
6.2.3 It is who we are, not what we see.
Our results show that reposting behaviour for previously unseen topics can be predicted by users’ profile and past behaviour, and is largely content-agnostic. In other words, who the users are plays a more central role than what information they are exposed to.
This aligns with recent LLM-agent simulations [55, 56] showing that echo chambers and extreme voices emerge irrespective of recommendation strategies. This observation at the population level echos our results on reposting prediction at the level of individual users. That is, if exchange of new information between individual users are largely content-agnostic, then the overall social dynamics in a population would be insensitive, or even immune, to recommendation algorithms.
Figure 2: In-distribution and out-of-distribution predictions by Decision Tree (DT) models using different types of features. User-related features U includes U-P, U-HA and U-HM.
7 Conclusion
Reposting prediction is a fundamental problem in modelling information diffusion. Existing works largely focused on textual data and only considered in-distribution evaluation. Our experimental results show that while message-based methods perform well in in-distribution scenarios, they show poor performance in out-of-distribution tasks when evaluated on repostings associated with previously unseen topics. We demonstrate that models incorporating user-related data always outperform message-based models. In particular, for out-of-distribution prediction, adding user-related data can significantly increase F1 score from 0.117 to 0.705.
Our findings suggest that a significant component of users’ reposting decisions, particularly for previously unseen topics, can be predicted based on users’ profile and past behaviour, and is largely content-agnostic.
We also argue that to ensure comprehensive evaluation while addressing practical needs, prediction works on social media should always include both in-distribution and out-of-distribution assessments.
Acknowledgments
The authors thank Ziwen Li, Berkem Billuroglu and Pei Lo for their contribution to the X data collection and preliminary analysis during their study at UCL. I.J.C. and V.L. would like to acknowledge support from the EPSRC grant EP/X031276/1.
Appendices
Appendix A Data Ethics
In our research, we collected data including messages and user profiles from publicly available sources (X platform, formerly Twitter). All data utilized in our study has been strictly limited to academic research purposes. The data has not been used to target or identify any individual personally. We have taken rigorous measures to ensure the anonymity and privacy of the data subjects. We have not disclosed any of the data to third parties.
Appendix B Hyper-parameters of Decision Tree Models
We searched for the optimal hyper-parameters of the DT models for each experiment setting. Searched values are provided in Table 6, where values marked with * are most frequently used. The parameter scale_pos_weight was only considered for imbalanced datasets (1:5 and 1:10).
Table 6: Hyper-parameters of Decision Tree models
Hyper-parameter Searched values
max_depth 6, 7, 8*, 9, 10
learning_rate 0.3*, 0.35, 0.4
n_estimators 100*, 150, 200
min_child_weight 1*, 2, 3
subsample 0.8, 0.9, 1.0*
scale_pos_weight [1*, 2, 3, 4, 5] (1:5 dataset)
[1.00*, 3.25, 5.50, 7.75, 10.00] (1:10 dataset)
Table 7: In-Distribution Prediction for Individual Hashtags using Temporal Split
Model DT-ALL DT-U DT-M
Dataset F1 Score (mean $\pm$ std)
[…]
Model NN-ALL NN-U NN-M BERT
Dataset F1 Score (mean $\pm$ std)
[…]
Appendix C Additional Temporal Prediction
Table 7 reports the additional temporal prediction results, i.e. in-distribution prediction results for individual hashtags using the temporal split. For each hashtag, we divide the dataset into 13 temporal windows with equal numbers of positive instances. The first 3 windows are used for training and the next for testing, rolling forward 1 window at each time step. The latest 10% of the training data is used for validation. Ten sequential predictions will be conducted. The results and conclusions are consistent with those presented in the main paper.
Appendix D Full Lists of Features
The following tables show the full lists of Message (M) features, User Profile (U-P) features, and User Historical Action (U-HA) features. The full list of User Historical Message (U-HM) features is in Table 1.
The following language models and tools are used for calculating message features.
  • Topic features: Twitter-roBERTa-base  [^4] https://huggingface.co/cardiffnlp, and Latent Dirichlet Allocation model from scikit-learn.
  • Language features: language-tool-python  [^5] https://github.com/jxmorris12/language_tool_python, TextBlob  [^6] https://github.com/sloria/textblob, and a Logistic Regression model trained on a public data source  [^7] https://www.kaggle.com/datasets/crowdflower/twitter-user-gender-classification.
  • Readability features: a Python implementation of several readability measures  [^8] https://github.com/andreasvc/readability/.
  • Sentiment features: VADER module from NLTK  [^9] https://www.nltk.org/index.html.
  • Emotion & Hate-Speech features: A Transformer-based library for SocialNLP tasks  [^10] https://github.com/pysentimiento/pysentimiento.
Table 8: Message (M) Features
Sub-type Feature Name Description
(# of features)
Topic (39) M_TopicM1 Likelihood that the message belongs to a specific topic in the 19-topic scheme 1.
M_TopicM19 (Same as above)
M_TopicMMain The most likely topic with the highest likelihood in the 19-topic scheme.
M_TopicMNum The number of identified topics with likelihood >> 0.5 in the 19-topic scheme.
M_TopicG1 Likelihood that the message belongs to a specific topic in the 6-topic scheme 2.
M_TopicG6 (Same as above)
M_TopicGMain The most likely topic with the highest likelihood in the 6-topic scheme.
M_TopicGNum The number of identified topics with likelihood >> 0.5 in the 6-topic scheme.
M_TopicLDA1 Likelihood that the message belongs to 1 of the 10 topics 3 identified by the LDA model.
M_TopicLDA10 (Same as above)
Language (10) M_CharNum The number of characters of the message.
M_WordNum The number of words of the message.
M_Grammar1 Scores for grammatical and spelling correctness on the level of words.
M_Grammar2 Scores for grammatical and spelling correctness of the whole message.
M_Polarity The orientation of the expressed tone of the message.
M_Subjectivity A measure of personal opinion and factual information contained in the message.
M_Irony A measure of how ironic the message is.
M_Offensive A measure of how offensive the message is.
M_Emoji The most likely emoji that describes the textual content.
M_Masculinity Whether the language style of the text is masculine.
Readability (11) M_Readability1 Different measures of how difficult the message in English is to understand.
(Readability measures includes: Kincaid, ARI, Coleman-Liau, FleschReadingEase,
M_Readability9         GunningFogIndex, SMOGIndex, LIX, RIX, DaleChallIndex)
M_Readability10 Count of words in the message that consists of three or more syllables.
M_Readability11 Count of words in the message that are not on the list used in DaleChallIndex.
Sentiment (5) M_Sentiment1 Negative sentiment score of the message.
M_Sentiment2 Neutral sentiment score of the message.
M_Sentiment3 Positive sentiment score of the message.
M_Sentiment4 Compound value of M_sentiment1, M_sentiment2, and M_sentiment3.
M_SentimentMain The most likely sentiment label according to M_sentiment4.
Emotion (8) M_Emotion1 The probability that the message expresses a specific type of emotion.
(Emotions studied: anger, joy, fear, disgust, surprise, sad, others.)
M_Emotion7 (Same as above)
M_EmotionMain The most likely emotion expressed by the message.
Hate-Speech (4) M_Hate1 A measure of aggressiveness of the message.
M_Hate2 A measure of hatefulness of the message.
M_Hate3 A measure of whether the message is targeting an individual (or entity).
M_HsNum The # of the above types of hate speech (above given thresholds) detected in the message.
Hashtag (1) M_hashtag The categorical label of the hashtag under study contained in the message.
1 The 19 topic-scheme includes: arts&culture, business&entrepreneurs, celebrity&pop_culture, diaries&daily_life, family, fashion&style, film_tv&video, fitness&health, food&dining, gaming, learning&educational, music, news&social_concern, other_hobbies, relationships, science&technology, sports, travel&adventure, youth&student_life. 2 The 6 topic-scheme includes: arts&culture, business&entrepreneurs, pop_culture, daily_life, sports&gaming, science&technology. 3 The top 3 keywords for each topic discovered by LDA model: 1) blm, amp, last; 2) nhs, covid, go; 3) john, mp, dido; 4) nhs, ios, one; 5) new, china, america; 6) loveisland, davide, think; 7) energy, brexit, energyprices; 8) avengers, marvel, news; 9) monkeypox, ukraine, russian; 10) covid, youtube, supercup.
Table 9: User Profile (U-P) Features
For Recipient (R)
Sub-type Feature Name Description
(# of features)
U-P_R_AccountAge The number of days since R’s account was registered (until 01/09/2022).
U-P_R_FollowerNum The number of R’s followers.
U-P_R_FolloweeNum The number of R’s followees (users that R is following).
U-P_R_TweetNum The number of total messages that R has posted since registration.
U-P_R_ListedNum The number of lists R is included in.
U-P_R_SpreadActivity The ratio of the # of R’s messages to the maximal # of messages among all users.
Profile U-P_R_FollowerNumDay The number of followers normalized by the age of this account.
(12) U-P_R_FolloweeNumDay The number of followees normalized by the age of R.
U-P_R_TweetNumDay The number of total messages normalized by the age of this account.
U-P_R_ListedNumDay The number of lists that the user is included in normalized by the age of this account.
U-P_R_ProfileVerified A binary label (0 or 1) indicating R is verified by X (1) or not (0).
U-P_R_ProfileUrl A binary label (0 or 1) indicating a URL is shown in R’s profile (1) or not (0).
U-P_R_LeaderRank A value to measure the influence of R. A variation of PageRank.
Network U-P_R_Indegree The number of users who have followed and interacted with R.
(3) U-P_R_FollowS A binary label (0 or 1) indicating R is following S (1) or not (0).
For Sender (S)
(Same as above except replacing ‘_R_’ with ‘_S_’ in feature names and replacing ‘R’ with ‘S’ in description.)
Table 10: User Historical Action (U-HA) Features
For Recipient (R)
Sub-type Feature Name Description
(# of features)
U-HA_R_TweetNum The number of R’s historical messages.
U-HA_R_TweetPercent Percentage of original-tweets/retweets/quotes/replies in R’s historical messages.
U-HA_R_RetweetPercent (See above)
Activity U-HA_R_QuotePercent (See above)
(7) U-HA_R_ReplyPercent (See above)
U-HA_R_InteractivePer Percentage of reposts in R’s historical messages.
U-HA_R_AverageInterval The average time interval (days) between the posting of R’s historical messages.
U-HA_R_RetweetedRate Average number of retweets/quotes/replies/likes received by R’s historical messages.
Popularity U-HA_R_QuotedRate (See above)
(4) U-HA_R_RepliedRate (See above)
U-HA_R_LikedRate (See above)
For Sender (S)
(Same as above except replacing ‘_R_’ with ‘_S_’ in feature names and replacing ‘R’ with ‘S’ in description.)
For both R and S
U-HA_RS_Mention The number of times that R mentioned S in R’s historical messages.
U-HA_RS_MentionPer Percentage of messages in which R mentioned S in R’s historical messages.
U-HA_SR_Mention The number of times that S mentioned R in S’s historical messages.
U-HA_SR_MentionPer Percentage of messages in which S mentioned R in S’s historical messages.
Interaction U-HA_RS_RepostLatency Latency (days) from the creation time of the message to the reposting.
(16) U-HA_SR_TORS1 A multi-dimensional vector to depict users’ relationship under different topics.
U-HA_SR_TORS10 (Same as above)
U-HA_SR_PathWidth The cosine distance between the message’s topic vector and the Topic-Oriented
Relationship Strength between S and R.