### v4

# Modelling the Spread of New Information on X

Abstract
There has been considerable interest in modelling the spread of information on X (formerly Twitter) using machine learning models. Here, we consider the problem of predicting the reposting of new information, i.e., when a user propagates information about a topic previously unseen by the user. In existing work, information and users are randomly assigned to a test or training set, ensuring that both sets are drawn from the same distribution. In the spread of new information, the problem becomes an out-of-distribution classification task.
Our experimental results reveal that while existing algorithms, which predominantly use features derived from the content of posts, perform well when the training and test distributions are the same, they perform much worse when the test set is out-of-distribution, i.e., when the topic of the testing data is absent from the training data. We then show that if the post features are supplemented or replaced with features derived from user profiles and past behaviours, the out-of-distribution prediction is greatly improved, with the F1 score increasing from 0.117 to 0.705. Our experimental results suggest that a significant component of reposting behaviour for previously unseen topics can be predicted from user profiles and past behaviours, and is largely content-agnostic.
Keywords Online Social Media $\cdot$ Social Content Sharing $\cdot$ Repost/Retweet $\cdot$ Out-of-Distribution Generalisation $\cdot$ Prediction $\cdot$ X (Twitter)
1 Introduction
Understanding and predicting user reposting behaviour on social networks is an area of interest across disciplines such as computer science, social science, political science, and marketing. This interest stems from the influence that individual behaviours have on information diffusion [40, 44, 39], including how information is spread during marketing [46] and political campaigns [47, 17, 13], and how misinformation and rumours spread [5, 8, 45].
The prediction of reposting has been formulated as a supervised classification task. The objective is to train a binary classifier to estimate the likelihood of a recipient reposting a post from a sender. In this setting, the dataset is typically divided into training and test sets through random assignment, ensuring that both sets are drawn from the same underlying data distribution for fair and stable evaluation.
On social media platforms, new topics appear regularly, e.g. those related to breaking news. Thus, in addition to the typical in-distribution classification, out-of-distribution prediction is also needed, where a model is trained on existing data related to previous topics (defined by hashtags) and is then used to predict a user’s reaction to a post related to a new, previously unseen, topic. Such a model can be useful for many applications, and it is more challenging, as it involves handling a distributional shift between user groups of known and unknown topics.
In this paper, we first review previous work on predicting repostings on X (formerly Twitter). Existing algorithms predominantly focused on the textual data, i.e., the text content of posts. We argue that features of the sender and the recipient are also relevant to the reposting. We study two types of data: (i) the post data, which contain the textual content of posts under consideration for reposting, (ii) the user data, which contain the sender and recipient’s profiles, their following lists, and their historical posts encoding their past activities, interests, and interactions with other users. We identified a total of 303 features from these data, including 78 post-related features and 225 user-related features. We consider three classification models: a decision tree model, a neural network model, and the BERT model [14], which is a widely studied Large Language Model (LLM). Using data collected from X, we conduct experiments under both in-distribution and out-of-distribution settings.
Our results show that while the prediction models, using only post-related features, perform well for in-distribution prediction, they perform much worse for out-of-distribution prediction. When the post-related features are supplemented or replaced with user-related features, the out-of-distribution performance greatly improves with the F1 score increasing from 0.117 to 0.705.
Our work highlights the importance of user-related features for reposting prediction—features that are currently underutilized. Interestingly, our experimental results suggest that a significant component of reposting behaviour for previously unseen topics is largely content-agnostic. This work provides novel insights into the mechanisms underlying information spreading on social networks.
2 Background
X (formerly Twitter) is a popular online social network. It allows users to publish short posts (formerly known as tweets), follow other users to curate their content feed (so called ‘timeline’), and interact with other users through actions, such as ‘repost’ and ‘like’. When a user creates a post, the post will be added to the timelines of the user’s followers as well as those who might be interested as determined by X’s recommendation algorithm.
When a user receives a post, it can ‘repost’ (formerly known as retweet) the post, ‘reply’ to the post by adding their response, or ‘quote’ the post by adding their comment. In our X data, reposts account for 78% of all posts, and replies and quotes account for 2%. In this study, we merge replies and quotes into reposts, because these actions add the post to the timelines of their followers and therefore have the same effect on information diffusion. A user can also ‘like’ a post to signal their endorsement, but this action does not add the post to the timelines of the user’s followers. The reposting function on X is similar to the ‘share’ function on Facebook. The ‘share’ function on Instagram and the ‘repost’ function on TikTok, however, are different as they forward posts privately to specific users or groups.
Source Type(#) Sub-type(#) Examples Note
Post-Related Features
Data(M) Text content of post M (78) Topic (39) M_TopicM5 Likelihood of a post related to topic ‘family’.
Language (10) M_Grammar1 Grammatical correctness of a post.
Readability (11) M_Readability1 A measure of difficulty in understanding a post.
Sentiment (5) M_Sentiment1 Negative sentiment score of a post.
Emotion (8) M_Emotion1 Probability of a post expressing emotion ‘anger’.
Hate-Speech (4) M_Hate1 A measure of aggressiveness of a post.
Hashtag (1) M_Hashtag The presence of a specific hashtag in a post.
User-Related Features for Sender $U_{S}$ and Recipient $U_{R}$
Data(U-P) U-P (30) Profile of $U_{S}$ (12) U-P_S_FollowerNum The number of followers of the Sender $U_{S}$.
Profile and friend list of $U_{S}$ and $U_{R}$ Profile of $U_{R}$ (12) U-P_R_FollowerNum The number of followers of the Recipient $U_{R}$.
U-P_S_FollowR A binary label indicating $U_{S}$ is following $U_{R}$ (1) or not (0).
U-P_R_FollowS A binary label indicating $U_{R}$ is following $U_{S}$ (1) or not (0).
Data(U-HA) Popularity of $U_{S}$ (4) U-HA_S_LikedRate Avg. # of ‘Likes’ received by latest 50 posts of $U_{S}$.
Public metrics of historical posts of $U_{S}$ and $U_{R}$ Popularity of $U_{R}$(4) U-HA_R_LikedRate Avg. # of ‘Likes’ received by latest 50 posts of $U_{R}$.
U-HA (38) Activity of $U_{S}$ (7) U-HA_S_InteractivePer Percentage of reposts in $U_{S}$’s latest 50 posts.
Activity of $U_{R}$(7) U-HA_R_InteractivePer Percentage of reposts in $U_{R}$’s latest 50 posts.
Interaction (16) U-HA_RS_Mention Number of $U_{R}$’s latest 50 posts that mentioned $U_{S}$.
Data(U-HM) U-HM (157) M features of $U_{S}$ (78) U-HM_S_Grammar1 (Each M feature is averaged over $U_{S}$’s latest 50 posts.)
Text content of historical posts of $U_{S}$ and $U_{R}$ M features of $U_{R}$ (78) U-HM_R_Grammar1 (Each M feature is averaged over $U_{R}$’s latest 50 posts.)
Topic Similarity (1) U-HM_SR_TopicSim Similarity between historical topics of 50 latest posts
of $U_{S}$ and $U_{R}$.
Table 1: Types and sub-types of features used in this study. Features are extracted from post data, Data(M), and user data, Data(U-P), Data(U-HA), and Data(U-HM).
2.1 Reposting Prediction
A reposting action involves three elements.
  • The content of a post ($M$). So far most works focus on the text content.
  • The sender ($U_{S}$), who created the post.
  • The recipient ($U_{R}$), i.e., the potential reposter, whose action is to be predicted.
In this study, we define reposting prediction as to predict whether a user (recipient) will repost a given post created by another user (sender), considering all of the three elements.
Specifically, we want to learn a binary classification model $f$, to predict the label $y$,
$f(M,U_{S},U_{R})\rightarrow y\in\{0,1\}\,$, (1)
where $y=1$ means the recipient will repost the post received from the sender, and $y=0$ otherwise.
2.2 Related Works
There are many related works on various aspects of reposting. Some works studied posts only, some focused on posts and their senders only, and some considered posts and their recipients only. For example, [36, 63] predicted the popularity of a post; [32, 52] predicted the incremental size of the spreading tree for a given post over time; [57] predicted the time gap between reposts of a given post; [49] predicted the probability of a given post being reposted (by any user), and [48, 26, 11, 58, 55] predicted the next user who is going to repost a given post based on the sequence of previous reposters of the post and the social relations among these reposters.
2.3 Prior Works on Reposting Prediction
Prior works on reposting prediction as defined in Equation (1) include a number of machine learning (ML) models, such as Support Vector Machines (SVM) [53, 34], Decision Trees  [53, 20, 65, 56], Logistic Regression  [59, 42], and factorisation models  [19, 27, 28, 22]. ML models are based on features extracted from raw data of posts and users. It is suggested in [53, 65] that the Decision Tree models often outperform other ML models.
Neural network (NN) models are also used for reposting prediction, including the Attention-based Convolutional Neural Network (SUA-ACNN) model [60] and the Bidirectional Long Short-Term Memory (Bi-LSTM) model [35]. NN models can learn directly from raw data and bypass the need for feature extraction.
Reposting prediction has also been studied from the perspective of recommender systems, which predict and rank the likelihood of candidate posts being reposted by a user [54, 19, 15, 1]. These studies used various prediction models, typically considered all the three elements, and evaluate their model performance at the ranked-list level.
2.4 Post Data vs User Data
Prior works on reposting prediction primarily relied on data of posts (including posts being reposted and the historial posts of users). For example, the ML models typically relied on features extracted from the textual content of post data, and the NN models often used the post data exclusively.
It is known, however, that user-related data, including users’ profile, their following list, and their past behaviour, interest and interactions, can have a significant influence on a user’s status, actions and decisions [34, 30, 21, 53, 61]. For example, [34] suggested that a recipient is more likely to repost a post from a sender if the recipient follows the sender, and the two users have shared interests and prior interactions. Prior works on recommender systems support similar observations  [9, 61, 51].
2.5 Out-of-Distribution Evaluation
Prior works on reposting prediction typically evaluated model performance by splitting the data into training and test sets using random sampling, where both sets contain data from the same set of topics [53, 34, 20, 65, 56, 59, 19, 27, 28, 22, 60, 35]. This setting, often referred to as in-distribution evaluation [3, 33], assumes that the training and test data are drawn from the same underlying distribution.
However, different topics may associate with user groups with distinct feature distributions. None of prior works explicitly evaluated generalisation across different topics in the out-of-distribution setting, that is to train a model using data of some topics (characterised by a set of hashtags) and then test the model using data of other topics, which can evaluate a model’s capacity in terms of predicting reposts associated with new, previously unseen topics.
3 Our Data for Reposting Prediction
3.1 X Data
We collected data from X using the former Twitter API with the Academic Access. We selected the 14 most popular hashtags in the trending list between July 27th and August 14th in 2022 and retrieved all available posts containing these hashtags during the same period around the world.
We obtained 63,336 posts, including 44,014 reposts (positive instances). These reposts satisfy two conditions: (1) each repost is created within 24 hours since its original post was posted; (2) the original post is also in our data. We set a 24-hour window because, according to the study of the half-life of X posts [41], users are unlikely to see a post more than 24 hours after its creation. Any reposting after 24 hours is likely influenced by external factors, such as TV or other social media platforms.
These posts (including the reposts) were created by 54,299 unique users, which include all senders and recipients of the reposts. For each user, we collected their profile, following list, and up to 50 latest historical posts. Work in [18, 4] suggest that 50 historical posts are sufficient to effectively characterise a user. A total of more than 2.6 million historical posts were collected. All users are anonymised in our data.
3.2 Post Data and Features
A post can have text and multimedia content, such as pictures, audio, and video. Same as in prior works, in this study, we only consider the text content of a post.
  • Data(M): the post data, contain the text content of the posts in our X data.
From the post data, we extract 78 post-related features, M, which are used in prior works for reposting prediction.
These features characterise a post from different perspectives. As shown in Table 1, we divide them into sub-types, including topic  [2, 23, 37, 65, 20, 22], language style [7, 19, 29, 12], readability [29], sentiment [25, 44, 62, 20], emotion [40, 44, 50, 20], hate speech  [38], and hashtag. These features are calculated using publicly available language models and tools (see Appendix). In addition, the hashtag feature is recorded as a categorical label [19, 20].
3.3 User Data and Features
In this study, the user data, Data(U), contain the following three types of user-related data.
  • Data(U-P): the profile of a user and the list of ‘friends’ that the user follows.
  • Data(U-HA): the metadata (or public metrics) of the latest 50 historical posts of a user, including (1) the type of a post, i.e., original post/repost/quote/reply, (2) the creation time of a post, (3) the number of reposts/quotes/replies/likes a post has received, and (4) any usernames mentioned in a post. These metadata encode useful information of a user’s previous behaviours, activity patterns, and its interactions with other users.
  • Data(U-HM): the textual content of the latest 50 historical posts of a user, which characterise a user’s historical interests and communication style.
As shown in Table 1, we can extract three types of user-related features from the above three types of user data, respectively. More details of the features are provided in the Appendix.
3.3.1 User Profile Features (U-P)
These features include the number of followers/followees of a user [40, 37, 64, 12], the number of lists that a user is included in [34, 53], the number of posts a user has created since its registration [34, 24, 37], whether a user is verified [19, 53], whether a URL is provided in a user’s profile [6], a user’s indegree [65] and LeaderRank [65], and whether a user is following another user [53].
3.3.2 User Historical Action Features (U-HA)
These features include, for example, the percentage of a user’s historical posts that are reposts [19, 65], the average time interval between a user’s posts [19], and the ratio of a sender’s historical posts that have mentioned a recipient [34, 53].
3.3.3 User Historical Post Features (U-HM)
These features include the aggregated post features (M) of a sender and those of a recipient, as well as an additional feature measuring topic similarity between the two users [65, 42].
3.4 Positive and Negative Instances
In this study, we define a positive instance as a reposting event $(M_{h},U_{S,t_{s}},U_{R,t_{r}})$, where the (original) post $M_{h}$ with a hashtag $h$ was created by a sender $U_{S}$ at $t_{s}$, and was reposted by a recipient $U_{R}$ at time $t_{r}$ (within 24 hours since $t_{s}$). Correspondingly, a negative instance is $(M^{*}_{h},U^{*}_{S,t_{s}},U_{R,t_{r}})$, where the same recipient $U_{R}$ did not repost another post $M^{*}_{h}$ with the same hashtag $h$ created by sender $U^{*}_{S}$ (where it is possible $U^{*}_{S}=U_{S}$). Here we consider negative instances that contained the same hashtag in the positive case and was created in the same 24-hour window, such that it is likely that $M^{*}_{h}$ was also exposed to the recipient and therefore makes the evaluation more relevant.
From the above X data, we created three datasets for reposting prediction with different ratios of positive to negative instances.
  • 1:1 dataset: For each positive instance, one negative instance is selected from our X data so that the post and the sender in the negative instance are most similar to those in the positive instance, i.e., the cosine distance between the feature vectors (containing all features of the post and the sender) of the positive instance and the chosen negative instance is the smallest among all negative instances.
  • 1:5 dataset: For each positive instance, the five most similar negative instances are selected, as defined above. This ratio is typically used in such experiments, such as in [60].
  • 1:10 dataset: For each positive instance, in addition to the 5 most similar negative instances, we add 5 negative instances randomly chosen from our X data.
Note that in all experiments in this study, if data is randomly split into training (and validation) and testing datasets, we remove any instance from the training dataset if the instance’s sender-recipient pair appears in the testing dataset. Such that, instances in the training and those in the testing datasets always have different sender-recipient pairs. This is to prevent information leakage and ensure a fair and rigorous evaluation.
4 Models for Reposting Prediction
In this study, we aim to examine the different contributions of post data and user data to the reposting prediction under both in-distribution and out-of-distribution settings. To focus on the aim, we studied three standard classification models to avoid complex tuning and optimisation associated with state-of-the-art models.
4.1 Decision Tree (DT) Models
The DT models remain a competitive choice for reposting prediction [65, 53]. In this study, we will examine how the performance of DT models is affected when different types of features are used. We use the XGBoost decision tree model. We searched for the optimal hyper-parameters for each data split of each dataset. The most frequently used hyper-parameters include the maximum depth of 8, the learning rate of 0.3, the estimators of 100, the minimum child weight of 1, the subsample ratio of 1.0, and the positive class weight of 1.0. For more details, please refer to the Appendix.
4.2 Neural Network (NN) Models
Figure 1: Neural network (NN) models based on different types of data.
The NN models have become a popular choice for reposting prediction [60, 35]. NN models can learn directly from raw data and bypass the need for feature extraction.
As shown in Figure 1, the NN models used in this study are extended from the SUA-ACNN model111 The SUA-ACNN model only used textual data, but it used textual data from not only the post being considered for reposting, Data(M), but also historical posts of the sender and the recipient of the reposting, Data(U-HM). In our study, we separate post data from user data.[60] by adding a component of Multi-Layer Perceptron (MLP)222 The MLP component has three dense layers with 128, 128 and 64 units, respectively. The activation function is set to ReLU. to incorporate additional user data, i.e., Data(U-P) and Data(U-HA). This allowed us to obtain a series of models using different types of data, including the NN-U model using the user data, the NN-M model using the post data only, and the NN-ALL model using all types of data. We set the hyperparameters as in [60]. NN models are trained with the Adam optimizer, the categorical cross-entropy loss, the learning rate of $0.001$, the batch size of $40$, and early stopping on validation loss.
4.3 BERT
We also study the BERT model [14], which is a widely studied Large Language Model (LLM). The model takes the post data as input and generates embeddings that represent the semantic meanings of the posts. The model then produced the prediction based on the embeddings. We used the recommended fine-tuning settings from [14]. We configure BERT-base with the AdamW optimizer, the learning rate of 2e52e-5, the batch size of $32$, and the dropout rate of $0.1$, with early stopping on validation loss. Cross-entropy loss was used as the objective function.
5 Experiments and Results
In this study, we conducted three experiments for reposting prediction using the above models. We examined whether the performance of reposting prediction, in terms of F1 score, can be improved by supplementing or replacing post data with user data. Firstly, we evaluated the models for in-distribution prediction on data of mixed hashtags; secondly, we compared the in-distribution prediction for individual hashtags; and thirdly, we conducted the out-of-distribution predictions.
See Appendix for definitions of the mean $\mu_{i}$, the standard deviation $\sigma_{i}$, the overall mean $\bar{\mu}$, and the overall standard deviation $\bar{\sigma}$.
5.1 Experiment I: In-Distribution Prediction For Mixed Hashtags
In this experiment, we mix data of all hashtags, and then performed Monte Carlo cross-validation 10 times, each randomly splitting the data 63/7/30 to the training, validation, and test sets.
In-Distribution Training on 70% randomly selected data, and testing on the remaining 30% data.
Decision Tree (DT) model Neural Network (NN) model LLM model
Model DT-ALL DT-U DT-M NN-ALL NN-U NN-M BERT
Feature(ALL) Feature(U) Feature(M) Data(All) Data(U) Data(M) Data(M)
Dataset F1 Score ($\mu$ $\pm$ $\sigma$)
[…]
Table 2: In-distribution prediction for mixed hashtags. Datasets are named by the ratio of positive and negative instances. Feature(ALL) includes both post-related features, Feature(M), and user-related features, Feature(U) which includes U-P, U-HA, and U-HM.
As shown in Table 2, the DT-ALL model using all types of features (303 in total) consistently shows significantly improved performance than other models using fewer features across datasets with different ratios of positive and negative instances. Among models with reduced feature sets, the DT-U model (utilizing user-related features) outperforms the DT-M model (utilizing post-related features).
The NN-ALL model utilizing all types of data produces the strongest performance among NN models. The NN-M model and the BERT model both utilize only post data and they achieve similar performance. Notably, these two models are outperformed by the NN-U model which used only user data.
These results consistently indicate that the inclusion of user data is valuable for reposting prediction.
As expected, the performance of all models degrades as the proportion of negative instances increases in the dataset, which aligns with the established understanding of classification challenges in imbalanced datasets. In the following of this paper, all results are based on the 1:5 dataset.
5.1.1 Comparison with Prior Works
In addition to the above results, we also conducted Experiment I using the decision tree model, TORS, in [65], which uses 37 features primarily derived from post data, and the neural network model, SUA-ACNN, in [60], which relies solely on post data. They achieved F1 scores of 0.737 and 0.739, respectively, which as expected are close to but lower than the DT-M model and the NN-M model.
Feature name Rank Feature Feature Import.
type sub-type value
U-P_R_FollowS 1 U-P network 0.229
U-HA_S_RetweetedRate 2 U-HA popularity 0.047
U-HA_S_LikedRate 3 U-HA popularity 0.045
U-HA_S_RepliedRate 4 U-HA popularity 0.042
U-HA_RS_MentionPer 5 U-HA interaction 0.015
M_TopicLDA7 1 6 M topic 0.013
U-HA_S_InteractivePer 7 U-HA activity 0.013
U-HA_S_TweetNum 8 U-HA activity 0.011
U-HA_S_RetweetPercent 9 U-HA activity 0.009
U-P_S_FollowR 10 U-P network 0.008
U-HA_RS_RepostLatency 11 U-HA interaction 0.008
U-HM_S_Readability7 12 U-HM readability 0.008
M_TopicLDA4 13 M topic 0.008
M_TopicLDA3 14 M topic 0.008
U-HM_S_TopicM7 15 U-HM topic 0.008
U-HM_S_TopicM11 16 U-HM topic 0.007
U-HM_S_TopicM19 17 U-HM topic 0.007
U-HA_RS_Mention 18 U-HA interaction 0.007
U-HA_S_TweetPercent 19 U-HA activity 0.007
U-P_S_Indegree 20 U-P network 0.007
# of features in each type: U-HA(10), U-HM(4), U-P(3), M(3).
Table 3: Top 20 most important features based on Decision Tree prediction. 1M_TopicLDA7 represents topics related to energy crisis, Brexit, and health issues. See full details of features in the Appendix.
5.1.2 Feature Importance
The feature importance value represents the relative contribution of a feature to the decision tree model’s predictive performance. Specifically, it is computed as the reduction in the loss function contributed by each feature when splitting nodes, with larger reductions indicating higher importance.
We obtained the feature importance values for all 303 features by training the DT-ALL model using all (instead of 70%) data in the 1:5 dataset. The feature importance values are normalised such that the sum for all features equals one. Our analysis showed that less than 1% of feature pairs has a Pearson coefficient value higher than 0.7 [16], which means most features contribute to the prediction.
Table 3 shows the top 20 most important features. It is notable that the majority of the top 20 features are user-related features, including 10 user historical action (U-HA) features, 4 user historical post (U-HM) features, and 3 user profile (U-P) features. The most important post (M) feature is merely ranked at the 6th position. This highlights the importance of user-related features for reposting predictions.
5.2 Experiment II: In-Distribution Prediction For Individual Hashtags
In this experiment, we conduct the in-distribution predictions for each of the 14 hashtags. We performed the same Monte Carlo cross-validation but on data of an individual hashtag.
Table 5 presents the in-distribution prediction results for each individual hashtag. The results are broadly consistent with those observed from the mixed-hashtag setting, with minor variations across individual hashtags. Specifically, the overall ranking of models remains the same: for decision trees, DT-ALL outperforms DT-U, which in turn outperforms DT-M; for neural network models, NN-ALL achieves the highest performance, followed by NN-U, with NN-M and BERT producing similar lowest performance. Both paired t-tests and Wilcoxon signed-rank tests confirm that the above performance rankings are statistically significant (with all values of p << 0.05). These findings indicate that the relative effectiveness of user and post data is stable across different in-distribution conditions.
The splitting schemes can also impact results. For the two splitting schemes, the mixed hashtag scope, which leverages data across all 14 hashtags, generally produces better performance than the individual hashtag scope. This suggests benefits from increased diversity and volume of the training data.
Note that, for rigorous evaluation, we have also conducted additional temporal experiments where data of each individual hashtag are divided into time windows according to the timestamp of repost events (see additional temporal experiments in Appendix). The results and conclusions remain the same: models using user data consistently outperform those using post data.
In-Distribution prediction For a given hashtag, training on 70% randomly selected data, test on remaining 30% data.
Decision Tree (DT) models Neural Network (NN) model LLM model
Model DT-ALL DT-U DT-M NN-ALL NN-U NN-M BERT
Feature(ALL) Feature(U) Feature(M) Data(All) Data(U) Data(M) Data(M)
Hashtag F1 Score ($\mu$ $\pm$ $\sigma$)
[…]
Table 4: In-distribution prediction for individual hashtags on the 1:5 dataset. See Appendix for the calcuation of $\mu$, $\sigma$, the overall mean $\bar{\mu}$, and the overall standard deviation $\bar{\sigma}$.
Out-of-distribution prediction For a given hashtag, training on data of the other 13 hashtags, test on the given hashtag.
Model DT-ALL DT-U DT-M NN-ALL NN-U NN-M BERT
Hashtag F1 Score ($\mu$ $\pm$ $\sigma$)
[…]
Table 5: Out-of-distribution prediction for individual hashtags on the 1:5 dataset. The expected F1 score for random guessing is 0.167, where positive and negative labels are randomly assigned at a ratio of 1:5.
5.3 Experiment III: Out-of-Distribution Prediction
We conducted out-of-distribution prediction for each of the 14 hashtags in our X data. For each given hashtag, we trained a model using data of the other 13 hashtags, and then tested the model with data of the given hashtag (the new, unseen topic). This split ensures an out-of-distribution condition, where the test posts come from a hashtag (topic) unseen during training 333 In our X data, we removed the small amount (less than 0.1%) of sender–recipient pairs that are overlapping across different hashtags. This ensures that sender–recipient pairs in the testing data are also unseen during training..
To obtain a stable and reliable estimate of model performance, for each given hashtag, we randomly divided the training data of the other 13 hashtags into 10 subsets, left one subset out for validation and used the other 9 subsets to train the model, and obtain prediction results. Then we repeated the process 10 times by leaving a different subset out each time. We evaluated the 10 trained models against the same test data of the given hashtag under study, and then obtained the mean and standard deviation of F1 scores.
Table 5 shows the out-of-distribution prediction by DT, NN and BERT models for each of the 14 hashtags on the 1:5 dataset. When based only on post data, all models show notably poor performance. Their overall F1 scores, $\bar{\mu}$, across 14 hashtags are as low as 0.117, 0.108, and 0.091, respectively, which are close to the expected F1 score of random guessing (0.167, see notes in Table 5). By comparison, the DT-ALL model and the NN-ALL model, using all types of data, perform significantly better.
Remarkably, the DT-U model and NN-U models, using only the user-related data, produce performance as good as, or even better than the DT-ALL and NN-ALL using all data. This suggests that user-related data are crucial for out-of-distribution prediction, where adding post-related data would have minimum benefit, or even negative impact, on performance. Both paired t-tests and Wilcoxon signed-rank tests confirm the performance rankings of DT-U >> DT-M, DT-ALL >> DT-M, and NN-U >> NN-ALL >> NN-M are statistically significant (with all values of p << 0.05).
The low performance of the NN-M model and the BERT model is unlikely due to any limitation in their architecture. Their underperformance is a result of exclusive reliance on post data. These two models learn from textual content of posts and therefore are bounded by the vocabularies present in the training data. This constrains their ability to generalise to posts of a new topic that are absent in the training data.
6 Discussion
6.1 In-Distribution vs. Out-of-Distribution
Figure 2 shows that the decision tree models’ performance for out-of-distribution prediction is always lower than that for in-distribution prediction. This is expected, as out-of-distribution prediction evaluates the model on a topic that is absent from the training set, whereas in-distribution prediction uses data of the same topics for both training and testing.
The gap between the two predictions is particularly pronounced for the model using post features M only, whose performance collapses for out-of-distribution prediction. This not only reveals the limited generalisation capability of models relying on post data, but also highlights the importance of conducting both in-distribution and out-of-distribution evaluations.
The model using U-HM features, extracted from the text content of users’ historical posts, performs similarly to the post-based model (M). Its out-of-distribution performance declines noticeably, though not as severely as M. This suggests that, beyond M, U-HM also contributes little to out-of-distribution prediction compared with features of user profiles (U-P) and user historical actions (U-HA).
In contrast, three user-based models (U, U-P, and U-HA) demonstrate significantly stronger predictive power, producing a performance as good as the model using ALL features for both in-distribution and out-of-distribution predictions. This is because these user-related features encode a user’s status and activity patterns that are independent from the latest trends on X and therefore can generalise well across different topics.
6.2 DT models vs. NN models
Performance of the DT-ALL model, using all types of features, is higher than that of the NN-ALL model, using all types of data, in both in-distribution experiments (for mixed and individual hashtags) and out-of-distribution experiments. This can perhaps be explained by the ability of decision trees to disregard uninformative features through their feature-splitting mechanism: at each node, the tree chooses the feature that best separates the classes, which not only reveals useful signals hidden in the data, but also effectively ignores weak or irrelevant features.
6.3 Users Shape Reposting Behaviour
Our results suggest that a significant component of reposting behaviour for previously unseen topics can be predicted from user profiles and past behaviours, and is largely content-agnostic. This supports the observation in [10] that cognitive and behavioural mechanisms are not peripheral details but fundamental determinants of information diffusion.
This, however, does not mean that reposting behaviour is independent of post content. Rather, our results suggest that both user-related and post-related factors contribute to reposting behaviour, as evidenced by the strong performance of the in-distribution predictions based on both user-related and post-related data.
The key difference is that content-related factors are inherently difficult to generalise to previously unseen topics due to semantic shifts; whereas user-related factors are largely content-agnostic, and therefore can maintain their predictive power when moving from in-distribution to out-of-distribution predictions.
6.4 Future Work
A recent work based on LLM-agent simulations [31, 43] showed that users can still form polarized echo chambers on social networks even without content-based recommendation algorithms. Future work should study how user characteristics, post content, and algorithmic exposure jointly influence information diffusion.
As multimodal content becomes increasingly common on social media, future work should study pictures and videos embedded in posts.
Figure 2: In-distribution and out-of-distribution predictions by Decision Tree (DT) models using different types of features, where U = U-P + U-HA + U-HM.
7 Conclusion
Reposting prediction is a fundamental problem in modelling information diffusion. Existing works largely focused on textual data and only considered in-distribution evaluation. Our experimental results show that while post-based methods perform well in in-distribution scenarios, they show poor performance in out-of-distribution tasks when evaluated on repostings associated with previously unseen topics. We demonstrate that models incorporating user-related data always outperform post-based models. In particular, for out-of-distribution prediction, adding user-related data can significantly increase the F1 score from 0.117 to 0.705.
Our findings suggest that a significant component of users’ reposting decisions, particularly for previously unseen topics, can be predicted based on user profiles and past behaviours, and is largely content-agnostic.
We also argue that to ensure comprehensive evaluation while addressing practical needs, prediction works on social networks should always include both in-distribution and out-of-distribution assessments.
Acknowledgments
The authors thank Ziwen Li, Berkem Billuroglu and Pei Lo for their contribution to the X data collection and preliminary analysis during their study at UCL. I.J.C. and V.L. would like to acknowledge support from the EPSRC grant EP/X031276/1.
Appendices
Appendix A Ethics Statement
In our research, we collected data including posts and user profiles from publicly available sources (X platform, formerly Twitter) under the platform’s terms of service. All data utilized in our study has been strictly limited to academic research purposes. The data has not been used to target or identify any individual personally. We have taken rigorous measures to ensure the anonymity and privacy of the data subjects. We have not disclosed any of the data to third parties.
Appendix B Computational Resources
All experiments in this study were conducted on a Linux-based workstation, with an AMD Ryzen 9 9950X CPU, 64 GB of system memory, and an NVIDIA GeForce RTX 4090 GPU. The operating system is Rocky Linux 9.6.
Our models are computationally efficient and compatible with standard GPU-equipped servers. No conclusions in this study depend on specialized hardware.
Appendix C Hyper-parameters of Decision Tree Models
We searched for the optimal hyper-parameters of the Decision Tree (DT) models for each experiment setting. The hyper-parameters were determined on the validation set. Searched values are provided in Table 6, where values marked with * are most frequently used. The parameter scale_pos_weight was only considered for imbalanced datasets (1:5 and 1:10).
Hyper-parameter Searched values
max_depth 6, 7, 8*, 9, 10
learning_rate 0.3*, 0.35, 0.4
n_estimators 100*, 150, 200
min_child_weight 1*, 2, 3
subsample 0.8, 0.9, 1.0*
scale_pos_weight [1*, 2, 3, 4, 5] (1:5 dataset)
[1.00*, 3.25, 5.50, 7.75, 10.00]
(1:10 dataset)
Table 6: Hyper-parameters of Decision Tree models
Model DT-ALL DT-U DT-M NN-ALL NN-U NN-M BERT
Dataset F1 Score (mean $\pm$ std)
[…]
Table 7: In-distribution prediction for individual hashtags using temporal split.
Appendix D Calculation of Mean and Standard Deviation
For each hashtag, we compute the mean $\mu_{i}$ and standard deviation $\sigma_{i}$ of the F1 scores across the 10 cross-validation folds. We then report the overall mean $\bar{\mu}$ and standard deviation $\bar{\sigma}$ across all 14 hashtags by treating the results from the 14 hashtags as equally weighted components of a mixture distribution.
$\bar{\mu}=\frac{1}{14}\sum_{i=1}^{14}\mu_{i}$ (2)
$\bar{\sigma}=\sqrt{\frac{1}{14}\sum_{i=1}^{14}\sigma_{i}^{2}+\frac{1}{14}\sum_{i=1}^{14}(\mu_{i}-\bar{\mu})^{2}}$ (3)
Appendix E Additional Results on a General 1:5 Dataset
In addition to the main experiments reported in the paper, where negative instances are constructed by selecting the five most similar non-reposted cases for each positive instance, we also conduct supplementary experiments on a more general 1:5 dataset, which serve as a robustness check and evaluate our models under a more general sampling strategy.
For convenience, we first define the set of posts in our raw data as $\mathcal{P}=\{(M_{h},U_{S,t_{s}})\}$ and the set of reposts $\mathcal{RP}=\{(M_{h},U_{S,t_{s}},U_{R,t_{r}})\}$ [^4] Note: Reposts whose original posts are absent from the raw data are excluded.. We further denote $\mathcal{U}=\{U_{t_{u}}\}$ as the set of users and their associated activity timestamps, which includes all $U_{S,t_{s}}$ from $\mathcal{P}$ and $U_{R,t_{r}}$ from $\mathcal{RP}$.
In the supplementary experiments, a negative instance $(M^{*}_{h},U^{*}_{S,t_{s}},U^{*}_{t_{u}})$ is generated by randomly pairing a post $(M^{*}_{h},U^{*}_{S,t_{s}})\in\mathcal{P}$ and a user $U^{*}_{t_{u}}\in\mathcal{U}$, on the condition that $(M^{*}_{h},U^{*}_{S,t_{s}},U^{*}_{t_{u}})\notin\mathcal{RP}$ and $t_{u}>t_{s}$. These conditions reflect the most fundamental causal requirements for a negative instance: a user $U^{*}_{t_{u}}$ can only repost a post $M^{*}_{h}$ if $U^{*}_{t_{u}}$ is active after $M^{*}_{h}$ is created, and no reposting event between $U^{*}_{t_{u}}$ and $M^{*}_{h}$ is observed in the data. Therefore, a negative instance is generated independently of a specific positive instance and without similarity constraints, which is more general than the strategy used in the main experiments.
The prediction results are shown in Table 8. The results and conclusions are consistent with those presented in the main paper.
Model DT-ALL DT-U DT-M
Experiment F1 Score
In-distribution (mixed) 0.916 0.907 0.775
In-distribution (individual) 0.901 0.894 0.740
Out-of-distribution 0.802 0.792 0.101
Table 8: Prediction results on the general 1:5 dataset. We focus on the Decision Tree (DT) models given their superior performance in the main experiments. Evaluations are conducted under three settings: (1) in-distribution (mixed) using data from all hashtags; (2) in-distribution (individual) using data from each individual hashtag, reported as the average across all individual hashtags; and (3) out-of-distribution, where the model is trained on 13 hashtags and tested on 1 remaining hashtags.
Appendix F Additional Results for Temporal Prediction
Table 7 reports the additional temporal prediction results, i.e., in-distribution prediction results for individual hashtags using the temporal split. For each hashtag, we divide the dataset into 13 temporal windows with equal numbers of positive instances. The first 3 windows are used for training and the next for testing, rolling forward 1 window at each time step. The latest 10% of the training data is used for validation. Ten sequential predictions will be conducted. The results and conclusions are consistent with those presented in the main paper.
Appendix G Full Lists of Features
The following tables show the full lists of Post (M) features, User Profile (U-P) features, and User Historical Action (U-HA) features. The full list of User Historical Post (U-HM) features is in Table 1.
The following language models and tools are used for calculating post features.
  • Topic features: Twitter-roBERTa-base 555 https://huggingface.co/cardiffnlp, and Latent Dirichlet Allocation model from scikit-learn.
  • Language features: language-tool-python 666 https://github.com/jxmorris12/language_tool_python, TextBlob 777 https://github.com/sloria/textblob, and a Logistic Regression model trained on a public data source 888 https://www.kaggle.com/datasets/crowdflower/twitter-user-gender-classification.
  • Readability features: a Python implementation of several readability measures 999 https://github.com/andreasvc/readability/.
  • Sentiment features: VADER module from NLTK 101010 https://www.nltk.org/index.html.
  • Emotion & Hate-Speech features: A Transformer-based library for SocialNLP tasks 111111 https://github.com/pysentimiento/pysentimiento.
Sub-type Feature Name Description
(# of features)
Topic (39) M_TopicM1 Likelihood that the post belongs to a specific topic in the 19-topic scheme 1.
M_TopicM19 (Same as above)
M_TopicMMain The most likely topic with the highest likelihood in the 19-topic scheme.
M_TopicMNum The number of identified topics with likelihood >> 0.5 in the 19-topic scheme.
M_TopicG1 Likelihood that the post belongs to a specific topic in the 6-topic scheme 2.
M_TopicG6 (Same as above)
M_TopicGMain The most likely topic with the highest likelihood in the 6-topic scheme.
M_TopicGNum The number of identified topics with likelihood >> 0.5 in the 6-topic scheme.
M_TopicLDA1 Likelihood that the post belongs to 1 of the 10 topics 3 identified by the LDA model.
M_TopicLDA10 (Same as above)
Language (10) M_CharNum The number of characters of the post.
M_WordNum The number of words of the post.
M_Grammar1 Scores for grammatical and spelling correctness on the level of words.
M_Grammar2 Scores for grammatical and spelling correctness of the whole post.
M_Polarity The orientation of the expressed tone of the post.
M_Subjectivity A measure of personal opinion and factual information contained in the post.
M_Irony A measure of how ironic the post is.
M_Offensive A measure of how offensive the post is.
M_Emoji The most likely emoji that describes the textual content.
M_Masculinity Whether the language style of the text is masculine.
Readability (11) M_Readability1 Different measures of how difficult the post in English is to understand.
(Readability measures includes: Kincaid, ARI, Coleman-Liau, FleschReadingEase,
M_Readability9 GunningFogIndex, SMOGIndex, LIX, RIX, DaleChallIndex)
M_Readability10 Count of words in the post that consists of three or more syllables.
M_Readability11 Count of words in the post that are not on the list used in DaleChallIndex.
Sentiment (5) M_Sentiment1 Negative sentiment score of the post.
M_Sentiment2 Neutral sentiment score of the post.
M_Sentiment3 Positive sentiment score of the post.
M_Sentiment4 Compound value of M_sentiment1, M_sentiment2, and M_sentiment3.
M_SentimentMain The most likely sentiment label according to M_sentiment4.
Emotion (8) M_Emotion1 The probability that the post expresses a specific type of emotion.
(Emotions studied: anger, joy, fear, disgust, surprise, sad, others.)
M_Emotion7 (Same as above)
M_EmotionMain The most likely emotion expressed by the post.
Hate-Speech (4) M_Hate1 A measure of aggressiveness of the post.
M_Hate2 A measure of hatefulness of the post.
M_Hate3 A measure of whether the post is targeting an individual (or entity).
M_HsNum The # of the above types of hate speech (above given thresholds) detected in the post.
Hashtag (1) M_hashtag The categorical label of the hashtag under study contained in the post.
Table 9: Post (M) Features. 1 The 19 topic-scheme includes: arts&culture, business&entrepreneurs, celebrity&pop_culture, diaries&daily_life, family, fashion&style, film_tv&video, fitness&health, food&dining, gaming, learning&educational, music, news&social_concern, other_hobbies, relationships, science&technology, sports, travel&adventure, youth&student_life. 2 The 6 topic-scheme includes: arts&culture, business&entrepreneurs, pop_culture, daily_life, sports&gaming, science&technology. 3 The top 3 keywords for each topic discovered by LDA model: 1) blm, amp, last; 2) nhs, covid, go; 3) john, mp, dido; 4) nhs, ios, one; 5) new, china, america; 6) loveisland, davide, think; 7) energy, brexit, energyprices; 8) avengers, marvel, news; 9) monkeypox, ukraine, russian; 10) covid, youtube, supercup.
For Recipient (R)
Sub-type Feature Name Description
(# of features)
U-P_R_AccountAge The number of days since R’s account was registered (until 01/09/2022).
U-P_R_FollowerNum The number of R’s followers.
U-P_R_FolloweeNum The number of R’s followees (users that R is following).
U-P_R_TweetNum The number of total posts that R has posted since registration.
U-P_R_ListedNum The number of lists R is included in.
U-P_R_SpreadActivity The ratio of the # of R’s posts to the maximal # of posts among all users.
Profile U-P_R_FollowerNumDay The number of followers normalized by the age of this account.
(12) U-P_R_FolloweeNumDay The number of followees normalized by the age of R.
U-P_R_TweetNumDay The number of total posts normalized by the age of this account.
U-P_R_ListedNumDay The number of lists that the user is included in normalized by the age of this account.
U-P_R_ProfileVerified A binary label (0 or 1) indicating R is verified by X (1) or not (0).
U-P_R_ProfileUrl A binary label (0 or 1) indicating a URL is shown in R’s profile (1) or not (0).
U-P_R_LeaderRank A value to measure the influence of R. A variation of PageRank.
Network U-P_R_Indegree The number of users who have followed and interacted with R.
(3) U-P_R_FollowS A binary label (0 or 1) indicating R is following S (1) or not (0).
For Sender (S)
(Same as above except replacing ‘_R_’ with ‘_S_’ in feature names and replacing ‘R’ with ‘S’ in description.)
Table 10: User Profile (U-P) Features
For Recipient (R)
Sub-type Feature Name Description
(# of features)
U-HA_R_TweetNum The number of R’s historical posts.
U-HA_R_TweetPercent Percentage of original-tweets/retweets/quotes/replies in R’s historical posts.
U-HA_R_RetweetPercent (See above)
Activity U-HA_R_QuotePercent (See above)
(7) U-HA_R_ReplyPercent (See above)
U-HA_R_InteractivePer Percentage of reposts in R’s historical posts.
U-HA_R_AverageInterval The average time interval (days) between the posting of R’s historical posts.
U-HA_R_RetweetedRate Average number of retweets/quotes/replies/likes received by R’s historical posts.
Popularity U-HA_R_QuotedRate (See above)
(4) U-HA_R_RepliedRate (See above)
U-HA_R_LikedRate (See above)
For Sender (S)
(Same as above except replacing ‘_R_’ with ‘_S_’ in feature names and replacing ‘R’ with ‘S’ in description.)
For both R and S
U-HA_RS_Mention The number of times that R mentioned S in R’s historical posts.
U-HA_RS_MentionPer Percentage of posts in which R mentioned S in R’s historical posts.
U-HA_SR_Mention The number of times that S mentioned R in S’s historical posts.
U-HA_SR_MentionPer Percentage of posts in which S mentioned R in S’s historical posts.
Interaction U-HA_RS_RepostLatency Latency (days) from the creation time of the post to the reposting.
(16) U-HA_SR_TORS1 A multi-dimensional vector to depict users’ relationship under different topics.
U-HA_SR_TORS10 (Same as above)
U-HA_SR_PathWidth The cosine distance between the post’s topic vector and the Topic-Oriented
Relationship Strength between S and R.
Table 11: User Historical Action (U-HA) Features